The measure of one angle is 5 times the measure of it's complement. What is the measure of each angle?
Answer by rfadrogane(214)

The measure of one angle is 5 times the measure of it's complement. What is the measure of each angle?
Sol'n:
let x = the 1st angle
90-x = the complement angle
x = 5(90-x)
x = 450-5x
6x = 450
x = 75 degrees ----answer
90-x = 15 degrees ----answer